what will be the change in speed of a body on application of force in 4seconds.If that force when exerted initially for 2 seconds raises the speed of the body from 4m/s to 8m/s

Grade:9

1 Answers

Manushri Manchanda
3 years ago
let force be F and mass be M F=M(V-U)/TF=M(8-4)/2F=2MNow when t=4 and f=2m2m=m(v-u)/48m/m=v-u8m/s=v-u